# GCL-OT: Graph Contrastive Learning with Optimal Transport for Heterophilic Text-Attributed Graphs

### v1

@@ Conclusion
This paper addresses the multi-granular heterophily challenges in TAGs. We propose GCL-OT, a contrastive learning framework with OT with tailored mechanisms to hierarchically align and fuse textual and structural representations. Specifically, for partial heterophily, GCL-OT employs the RealSoftMax operator to identify important structure-text relations. For complete heterophily, GCL-OT employs a filter prompt to distinguish between alignable and non-alignable embeddings in the transport-based contrastive loss. For latent homophily, GCL-OT leverages OT assignments as extra supervision in the contrastive loss to capture hidden homophilic patterns. Theoretical analysis and extensive experiments on both homophilic and heterophilic settings can demonstrate the effectiveness and robustness of GCL-OT.
@@ Acknowledgments
Optimal transport (OT) (Monge1781; Villani2009OT)is a mathematical framework for measuring distances between distributions, finding the most cost-efficient way to transform one distribution into another (lin2023multi; Vincent2022OTGNN; Xu2023GWF). In Graph Contrastive Learning (GCL), traditional InfoNCE loss usually relies on hard alignment and explicit positive and negative samples, which may lead to representation bias. To this end, researchers begin to use OT to introduce a soft alignment mechanism (wang2023galopa). Several methods employ OT to enhance node representation learning, which subsequently benefits downstream tasks such as node clustering  (Zhang2024CSOT; Wang2024PVCCL; deng2025THESAURUS) and node classification (xie2024SGEC; Amadou2025FOSSIL; Zhu2022RoSA).
The node classification method aims to solve the robustness problems caused by non-aligned contrast views and structural differences in the graph by introducing OT distance for soft alignment, subgraph generation, or cross-view comparison, thereby improving the accuracy and stability of node classification (deng2025THESAURUS; deng2025THESAURUS). For example, in order to alleviate the uniform effect and cluster assimilation problems, THESAURUS (deng2025THESAURUS) introduces learnable semantic prototype graphs, employs Gromov–Wasserstein OT to align node embeddings with these prototypes, and uses cross-view prediction of cluster assignment as its learning objective. Considering that existing GCL methods often struggle to exploit structural patterns and node similarities fully, FOSSIL (Amadou2025FOSSIL) introduces an adaptive subgraph generator and formulates a subgraph-level contrastive loss using the Fused Gromov-Wasserstein distance as the similarity measure. Beyond clustering and classification, OT has also shown promise in graph few-shot learning (Liu2025STAR) and node anomaly detection (Wang2023ACT).
@@
  • GCN (kipf2017GCN): A spectral‑based graph convolutional network that aggregates one‑hop neighborhood features for semi‑supervised node classification.
  • GAT (velivckovic2018GAT): Adds multi‑head attention in message passing to adaptively weight neighbor contributions.
  • GraphSAGE (hamilton2017SAGE): Uses learned aggregate functions over sampled neighborhoods, supports inductive learning on large graphs.
  • GIANT‑BERT (chien2022giant): Learns node textual features by self-supervised neighbor prediction using BERT. These GIANT-generated features are then fed into a GNN classifier.
  • GLEM‑DeBERTa (zhao2023GLEM): Trains a language model and GNN iteratively under a variational EM framework, the DeBERTa variant is used to generate text-informed node features that the GNN consumes.
  • ENGINE‑LLAMA (zhu2024ENGINE): Introduces an auxiliary, parameterized LM side‑structure to inject graph structure into LLaMA during fine‑tuning, combining graph-aware features with efficient inference and early exit strategies.
  • TAPE‑GCN / TAPE‑RevGAT / TAPE‑SAGE (he2024TAPE): First generate node explanations via LLM prompting, those explanations are translated into features that feed into GNN classifiers based on GCN, RevGAT, and GraphSAGE backbones respectively, achieving strong performance via enriched semantics.
  • SimTeG‑e5 (duan2023simteg): A frustratingly simple two-stage method: fine-tune a text embedding model (e.g., E5) on node classification, and use those embeddings to train a GNN—demonstrating that improved text embeddings alone substantially boost GNN performance.
  • LEMP‑TAPE (wang2025LEMP4HG): LEMP‑TAPE extends TAPE by using LEMP4HG’s LM‑enhanced message passing: for selected edges (particularly in heterophilic graph regions), a language model produces a connection analysis given paired node texts, that analysis is encoded and fused via gating with source and target node embeddings to synthesize enriched messages. Only selected edges are enhanced via active learning heuristic (MVRD), reducing overhead while improving robustness on heterophilic graphs.
  • DGI (velickovic2019dgi): An unsupervised graph representation learning method that maximizes mutual information between local patch representations and a global graph summary using a GCN encoder. It achieves strong performance on node classification without task-specific labels.
  • GRACE (zhu2020GRACE): Learns node embeddings by contrasting augmented views of the graph. GRACE improves notably over DGI, especially on datasets with sparse node features, by exploiting topology-aware negative sampling and graph augmentations.
  • ConGraT (brannon2024congrat): A self‑supervised method jointly pretraining a language model and a GNN via contrastive alignment of text and node representations in the same latent space, balancing textual and structural signals for downstream tasks.
@@
Thus we obtain that the upper bound for $\mathcal{L}_{\text{MHA}}$ has the form of $\mathbb{E}_{i}\left[\log\frac{\exp(s_{ii}/\tau)}{\sum_{j=1}^{N}\exp(s_{ij}/\tau)}+\log\frac{\exp(s_{ii}/\tau)}{\sum_{j=1}^{N}\exp(s_{ji}/\tau)}\right]$, which is the same as the objective in the CLIP-style symmetric InfoNCE, i.e., the sum of the two directions (oord2018representation; poole2019variational; radford2021learning; zhu2020GRACE; zhu2025GraphCLIP).
@@
We formally illustrate all metrics used in our empirical analysis. Following (Pei2020GeomGCN), the node-level homophily $\mathcal{H}_{N}$, which measures the average proportion of neighbors sharing the same label, is calculated as
@@
Following (zhu2020H2GCN), the edge-level homophily $\mathcal{H}_{E}$, which denotes the proportion of edges connecting nodes with the same label, is calculated as
@@
Inspired by (luan2023CSBMPBE), the proportion of nodes whose neighbors have mixed labels $\mathcal{R}_{\text{NYM}}$ is calculated as
@@
where $\tau$ is typically set to 0.5 following (wang2024understanding).
@@
Table 4 outlines the prompts used for various datasets. Following TAPE (he2024TAPE), each prompt includes the content and title of the text, followed by a task-specific question. The question is formulated to query the model about a particular aspect of the text and request an explanation for the prediction. The answer section is left blank for the model to fill in. Generally, our analysis finds that the current instructions allow the LLM to produce output that conforms well to the expected format without significant deviations, allowing the answers to be straightforwardly extracted from the text output of the LLM.
@@
Dataset Node Edge Feature Class $H_{N}$ $H_{N}$ $H_{E}$ $R_{NYS}$ $R_{NYD}$ $R_{NYM}$ $R_{UNS}$ $R_{NWD}$ $R_{NTD}$ $R_{UTS}$
Cora 2,708 5,278 1,433 7 0.810 82.52 81.00 65.58 5.80 28.62 99.17 85.77 18.22 82.98
Pubmed 19,717 44,324 500 3 0.802 79.24 80.20 64.43 12.79 22.78 99.95 81.12 13.71 94.80
Products 54,025 74,420 100 47 0.808 78.97 80.80 70.11 16.21 13.68 99.95 94.89 29.72 54.19
arXiv 169,343 1,166,243 128 40 0.655 63.53 65.50 28.35 12.28 59.36 99.93 80.09 23.59 19.25
Amazon-ratings 24,492 186,100 300 5 0.380 37.57 38.00 2.55 16.08 81.37 99.95 76.70 17.06 85.53
arXiv23 46,198 78,548 128 40 0.647 29.66 64.70 54.15 19.31 26.53 99.98 83.87 17.93 51.34
Wisconsin 265 530 1,703 5 0.200 18.68 21.90 4.53 60.38 35.09 98.83 88.06 16.59 95.67
Cornell 195 304 1,703 5 0.132 12.12 13.20 4.62 74.36 21.03 99.06 89.54 18.87 96.08
Texas 187 328 1,703 5 0.120 10.68 10.80 3.74 77.54 18.72 99.56 89.52 18.88 97.63
Actor 4,416 12,172 932 5 0.560 60.86 56.00 34.40 16.19 49.41 99.76 72.74 23.75 77.57
@@
  • Cora, Pubmed (yang2016revisiting; sen2008collective_pubmed), arXiv (hu2020ogb), and arXiv23 (he2024TAPE) are four citation networks that are considered classic homophilic graphs. In these graphs, nodes represent papers, edges represent citation relationships between two papers, and labels indicate the research topic of each paper.
  • Products (hu2020ogb) is a co-purcharse network. In this graph, nodes represent products, edges represent co-purchase relationships between two products, and labels indicate the product category of each product.
  • Cornell, Texas, and Wisconsin (Pei2020GeomGCN) are derived from computer science department websites where webpages serve as nodes, hyperlinks as edges, and entity types as labels.
  • Actor (tang2009social) is an actor co-occurrence network with nodes representing actors, edges denoting their collaborations, and labels indicating the Wikipedia-derived actor categories of each actor.
  • Amazon-ratings (platonov2023critical) is a product co-purchasing network where nodes represent products, edges link products frequently bought together, and labels correspond to rating levels assigned to each product.

### v2

@@ Conclusion
This paper addresses the multi-granular heterophily challenges in TAGs. We propose GCL-OT, a contrastive learning framework with OT with tailored mechanisms to hierarchically align and fuse textual and structural representations. Specifically, for partial heterophily, GCL-OT employs the RealSoftMax operator to identify important structure-text relations. For complete heterophily, GCL-OT employs a filter prompt to distinguish between alignable and non-alignable embeddings in the transport-based contrastive loss. For latent homophily, GCL-OT leverages OT assignments as extra supervision to capture hidden homophilic patterns. Theoretical analysis and extensive experiments in both homophilic and heterophilic settings can demonstrate the effectiveness and robustness of GCL-OT.
@@ Acknowledgments
Optimal transport (OT) (Monge 1781; Villani 2009)is a mathematical framework for measuring distances between distributions, finding the most cost-efficient way to transform one distribution into another (Lin et al. 2023; Vincent-Cuaz et al. 2022; Xu et al. 2023). In Graph Contrastive Learning (GCL), traditional InfoNCE loss usually relies on hard alignment and explicit positive and negative samples, which may lead to representation bias. To this end, researchers begin to use OT to introduce a soft alignment mechanism (Wang et al. 2023b). Several methods employ OT to enhance node representation learning, which subsequently benefits downstream tasks such as node clustering  (Zhang et al. 2024b; Wang et al. 2024c; Deng et al. 2025) and node classification (Xie and Giraldo 2024; SANGARE et al. 2025; Zhu et al. 2022).
The node classification method aims to solve the robustness problems caused by non-aligned contrast views and structural differences in the graph by introducing OT distance for soft alignment, subgraph generation, or cross-view comparison, thereby improving the accuracy and stability of node classification (Deng et al. 2025, 2025). For example, in order to alleviate the uniform effect and cluster assimilation problems, THESAURUS (Deng et al. 2025) introduces learnable semantic prototype graphs, employs Gromov–Wasserstein OT to align node embeddings with these prototypes, and uses cross-view prediction of cluster assignment as its learning objective. Considering that existing GCL methods often struggle to exploit structural patterns and node similarities fully, FOSSIL (SANGARE et al. 2025) introduces an adaptive subgraph generator and formulates a subgraph-level contrastive loss using the Fused Gromov-Wasserstein distance as the similarity measure. Beyond clustering and classification, OT has also shown promise in graph few-shot learning (Liu et al. 2025b) and node anomaly detection (Wang et al. 2023a).
@@
  • GCN (Kipf and Welling 2017): A spectral‑based graph convolutional network that aggregates one‑hop neighborhood features for semi‑supervised node classification.
  • GAT (Veličković et al. 2018): Adds multi‑head attention in message passing to adaptively weight neighbor contributions.
  • GraphSAGE (Hamilton et al. 2017): Uses learned aggregate functions over sampled neighborhoods, supports inductive learning on large graphs.
  • GIANT‑BERT (Chien et al. 2022): Learns node textual features by self-supervised neighbor prediction using BERT. These GIANT-generated features are then fed into a GNN classifier.
  • GLEM‑DeBERTa (Zhao et al. 2023): Trains a language model and GNN iteratively under a variational EM framework, the DeBERTa variant is used to generate text-informed node features that the GNN consumes.
  • ENGINE‑LLAMA (Zhu et al. 2024): Introduces an auxiliary, parameterized LM side‑structure to inject graph structure into LLaMA during fine‑tuning, combining graph-aware features with efficient inference and early exit strategies.
  • TAPE‑GCN / TAPE‑RevGAT / TAPE‑SAGE (He et al. 2024): First generate node explanations via LLM prompting, those explanations are translated into features that feed into GNN classifiers based on GCN, RevGAT, and GraphSAGE backbones respectively, achieving strong performance via enriched semantics.
  • SimTeG‑e5 (Duan et al. 2023): A frustratingly simple two-stage method: fine-tune a text embedding model (e.g., E5) on node classification, and use those embeddings to train a GNN—demonstrating that improved text embeddings alone substantially boost GNN performance.
  • LEMP‑TAPE (Wang and Cheng 2025): LEMP‑TAPE extends TAPE by using LEMP4HG’s LM‑enhanced message passing: for selected edges (particularly in heterophilic graph regions), a language model produces a connection analysis given paired node texts, that analysis is encoded and fused via gating with source and target node embeddings to synthesize enriched messages. Only selected edges are enhanced via active learning heuristic (MVRD), reducing overhead while improving robustness on heterophilic graphs.
  • DGI (Veličković et al. 2019): An unsupervised graph representation learning method that maximizes mutual information between local patch representations and a global graph summary using a GCN encoder. It achieves strong performance on node classification without task-specific labels.
  • GRACE (Zhu et al. 2020b): Learns node embeddings by contrasting augmented views of the graph. GRACE improves notably over DGI, especially on datasets with sparse node features, by exploiting topology-aware negative sampling and graph augmentations.
  • ConGraT (Brannon et al. 2024): A self‑supervised method jointly pretraining a language model and a GNN via contrastive alignment of text and node representations in the same latent space, balancing textual and structural signals for downstream tasks.
@@
Thus we obtain that the upper bound for $\mathcal{L}_{\text{MHA}}$ has the form of $\mathbb{E}_{i}\left[\log\frac{\exp(s_{ii}/\tau)}{\sum_{j=1}^{N}\exp(s_{ij}/\tau)}+\log\frac{\exp(s_{ii}/\tau)}{\sum_{j=1}^{N}\exp(s_{ji}/\tau)}\right]$, which is the same as the objective in the CLIP-style symmetric InfoNCE, i.e., the sum of the two directions (Oord et al. 2018; Poole et al. 2019; Radford et al. 2021; Zhu et al. 2020b, 2025).
@@
We formally illustrate all metrics used in our empirical analysis. Following (Pei et al. 2020), the node-level homophily $\mathcal{H}_{N}$, which measures the average proportion of neighbors sharing the same label, is calculated as
@@
Following (Zhu et al. 2020a), the edge-level homophily $\mathcal{H}_{E}$, which denotes the proportion of edges connecting nodes with the same label, is calculated as
@@
Inspired by (Luan et al. 2023), the proportion of nodes whose neighbors have mixed labels $\mathcal{R}_{\text{NYM}}$ is calculated as
@@
where $\tau$ is typically set to 0.5 following (Wang et al. 2024b).
@@
Table 4 outlines the prompts used for various datasets. Following TAPE (He et al. 2024), each prompt includes the content and title of the text, followed by a task-specific question. The question is formulated to query the model about a particular aspect of the text and request an explanation for the prediction. The answer section is left blank for the model to fill in. Generally, our analysis finds that the current instructions allow the LLM to produce output that conforms well to the expected format without significant deviations, allowing the answers to be straightforwardly extracted from the text output of the LLM.
@@
Dataset Node Edge Feature Class $H_{N}$ $H_{E}$ $R_{NYS}$ $R_{NYD}$ $R_{NYM}$ $R_{UNS}$ $R_{NWD}$ $R_{NTD}$ $R_{UTS}$
Cora 2,708 5,278 1,433 7 82.52 81.00 65.58 5.80 28.62 99.17 85.77 18.22 82.98
Pubmed 19,717 44,324 500 3 79.24 80.20 64.43 12.79 22.78 99.95 81.12 13.71 94.80
Products 54,025 74,420 100 47 78.97 80.80 70.11 16.21 13.68 99.95 94.89 29.72 54.19
arXiv 169,343 1,166,243 128 40 63.53 65.50 28.35 12.28 59.36 99.93 80.09 23.59 19.25
Amazon-ratings 24,492 186,100 300 5 37.57 38.00 2.55 16.08 81.37 99.95 76.70 17.06 85.53
arXiv23 46,198 78,548 128 40 29.66 64.70 54.15 19.31 26.53 99.98 83.87 17.93 51.34
Wisconsin 265 530 1,703 5 18.68 21.90 4.53 60.38 35.09 98.83 88.06 16.59 95.67
Cornell 195 304 1,703 5 12.12 13.20 4.62 74.36 21.03 99.06 89.54 18.87 96.08
Texas 187 328 1,703 5 10.68 10.80 3.74 77.54 18.72 99.56 89.52 18.88 97.63
Actor 4,416 12,172 932 5 60.86 56.00 34.40 16.19 49.41 99.76 72.74 23.75 77.57
@@
  • Cora, Pubmed (Yang et al. 2016; Sen et al. 2008), arXiv (Hu et al. 2020), and arXiv23 (He et al. 2024) are four citation networks that are considered classic homophilic graphs. In these graphs, nodes represent papers, edges represent citation relationships between two papers, and labels indicate the research topic of each paper.
  • Products (Hu et al. 2020) is a co-purcharse network. In this graph, nodes represent products, edges represent co-purchase relationships between two products, and labels indicate the product category of each product.
  • Cornell, Texas, and Wisconsin (Pei et al. 2020) are derived from computer science department websites where webpages serve as nodes, hyperlinks as edges, and entity types as labels.
  • Actor (Tang et al. 2009) is an actor co-occurrence network with nodes representing actors, edges denoting their collaborations, and labels indicating the Wikipedia-derived actor categories of each actor.
  • Amazon-ratings (Platonov et al. 2023) is a product co-purchasing network where nodes represent products, edges link products frequently bought together, and labels correspond to rating levels assigned to each product.